The displacement of a particle is given as x = 20 Cos (200 - ∅). Calculate the maximum acceleration of the particle
1
Expert's answer
2021-03-08T08:29:53-0500

"x=20\cos(200t-\phi)"


"a=d^2x/dt^2=-20\cdot200^2\cdot\cos(200t-\phi)"


"a_{max}=-20\cdot200^2=-8\cdot10^5\ (m/s^2)" . Answer